Project K costs $75,000, its expected cash inflows are $13,000 per year for 8 years, and its WACC is 12%. What is the project's MIRR? Round your answer to two decimal places.
Answer #1
FV of cash inflows = 13000*(1.12^8-1)/0.12 = $        1,59,896.01
MIRR = (159896.01/75000)^(1/8)-1 = 9.93%